A motorist covered a journey of 690km in 8 hours. He covered part of the journey at 90km/h and the rest at 80km/h. What was the distance that he covered at 80km/h?
Answer by richwmiller(17219)

a+b=8,
90*a+80*b=690
a=8-b
90*(8-b)+80*b=690
720-90b+80b=690
30=10b
b=3 hours at 80 mph
a=5 hours at 90 mph
check
90*5+80*3=690
450+240=690
690=690
ok